Dr. John Paul Stapp was U.S. Air Force officer who studied the effects of extreme deceleration on the human body. On December 10, 1954, Stapp rode a rocket sled, accelerating from rest to a top speed of 282 m/s (1015 km/h) in 5.00 s, and was brought jarringly back to rest in only 1.40 s!


1
Expert's answer
2021-02-25T11:23:30-0500

Given,

Initial velocity "(v_i)=0"

Final velocity "(v_f)=282 m\/s"

Time taken "(t)=5.0 sec"

Acceleration of the particle "(a) = \frac{v_f-v_i}{t}"

Now, substituting the values,

"a=\frac{282}{5} m/s^2"

"\Rightarrow a = 56.4 m/s^2"

Acceleration in terms of g ,

"a_g= \frac{56.4g}{9.8}=5g m/s^2"

It comes in rest in 1.4s so retardation "a_r= \frac{-282}{1.4} m/s^2"

"=201.43 m\/s^2"

Retardation in terms of g,

"a_{rg}=\frac{201.43g}{9.8} m/s^2"


"=-20.5g m\/s^2"